OpenStudy (anonymous):

Solve for x. 8 – 3x – x = 12 A. –10 B. –1 C. –2 D. –5

pooja195 (pooja195):

combine like terms -3x-x=?

OpenStudy (anonymous):

hmmmm?

pooja195 (pooja195):

-3 -1 ?

OpenStudy (anonymous):

oh i see thanks

OpenStudy (pawanyadav):

8-4x=12 -4x=12-8 -4x=4 X =-1